OpenStudy (anonymous):

Tickets for a concert cost $20 for a kid and $30 for a parent If 500 tickets were sold for a total of $12000 what was the ratio of the number of kid tickets sold to the # of parent tickets?

OpenStudy (anonymous):

So if you have x amount of tickets sold to kids y amount of tickets sold to adults and we know that the total number of tickets was 500 so x+y=500 Then for every kid ticket sold, it was $20 so total is 20*x every adult ticket was $30, so total is 30*y we know they make $12000 so 20x+30y=12000 so we have two equations to work with x+y=500 20x+30y+12000 if we scale up we can see that 20x+20y=10000 no eliminate '20x' so 30y-20y=12000-10000 so 10y=2000 y=200 ie number of adult tickets sold if x+y=500 and y=200 then x=300 so the ratio is 300:200 or 3:2 Hope that helps!

OpenStudy (anonymous):

THANK YOU <3

OpenStudy (anonymous):

no problem :)